OpenStudy (anonymous):

simplify the following quotient of complex numbers into the form a+bi i/8+8i

OpenStudy (lgbasallote):

\[\frac{i}{8 + 8i} \implies \frac{i}{8+8i} \times \frac{8 - 8i}{8-8i}\] does that help?

OpenStudy (anonymous):

i'm not sure i understand the formula?

OpenStudy (lgbasallote):

it's not a formula..it's multiplying by the conjugate`

OpenStudy (lgbasallote):

you are familiar with that arent you?

OpenStudy (anonymous):

yes kind of..... is that the opposite?

OpenStudy (lgbasallote):

it has the opposite sign of the denominator your denominator is 8 + 8i therefore you multiply numerator and denominator by 8-8i

OpenStudy (anonymous):

ok ty

OpenStudy (lgbasallote):

welcome